OpenStudy (anonymous):

What is 3 X 2 5/6 = ? Help Please and thank you :D

OpenStudy (anonymous):

is the x suppose to be a variable?

OpenStudy (anonymous):

if it's multiplication, I think you just have to multiply 3 and 2 and then add 5/6 so it's 6 5/6

OpenStudy (anonymous):

no it's a multiplication sign

OpenStudy (anonymous):

I know the answer but I just don't know how to do it. the answer is 8 3/6 or 8 1/2.

OpenStudy (hoblos):

3 x 2 5/6 = 3(2+ 5/6) = 6 + 15/6 = 8 3/6

OpenStudy (anonymous):

thank you hoblos